OpenStudy (anonymous):

Four coins are tossed at the same time. What is the probability that exactly 3 heads will appear? 1/16 1/8 1/32 1/4

OpenStudy (perl):

you want P(HHH)

OpenStudy (perl):

P(HHH) = P(H)*P(H)*P(H) = 1/2 * 1/2 * 1/2

OpenStudy (anonymous):

=0.215

OpenStudy (perl):

.125 i think

OpenStudy (anonymous):

oh ok

OpenStudy (perl):

but your choices are given as fractions

OpenStudy (anonymous):

is the answer 1/8?